OpenStudy (anonymous):

Find a quadratic equation with roots -1 + 4i and -1 - 4i. Need help please!

OpenStudy (midhun.madhu1987):

Sum of roots = -1 + 4i -1 - 4i = -2 Product of roots = (-1 + 4i)(-1 - 4i) = 1 + 16 = 17

OpenStudy (midhun.madhu1987):

 this is the equation of a quadratic equation